OpenStudy (anonymous):

Evaluate 64 to the two thirds power. a 16 b 4 c 8 d 12

OpenStudy (welshfella):

This is the square of the cube root of 64 Find the cube root first

OpenStudy (welshfella):

The top of the fraction gives the power and the denominator gives the root.

OpenStudy (anonymous):

so which answer is it you think?

OpenStudy (welshfella):

Well what is the cube root of 64? What number multiplied by itself 3 times gives 64 An example of cube. Root is 2 is the cube rot of 8 because 2 * 2. * 2 = 8

OpenStudy (anonymous):

??

OpenStudy (anonymous):

which letter would that be ? a b or c?

OpenStudy (welshfella):

Also cube root of 27 is 3 because 27 = 3 * 3 * 3 = 27

OpenStudy (welshfella):

I have just given 2 examples of cube roots to help you find the cube root of 64

OpenStudy (anonymous):

okay i don't want math help i just want the answer please and thanks

OpenStudy (welshfella):

Or use your calculator to find it

OpenStudy (anonymous):

im in a really hard online class I'm just trying to pass

OpenStudy (welshfella):

4 times 4 times 4 = what?

OpenStudy (anonymous):

okay thanks you could've just said the answer is a 16 instead of wasting like 20 min

OpenStudy (welshfella):

It is not wasted time if I help you understand how to do the problem

OpenStudy (anonymous):

nah you didn't help me

OpenStudy (welshfella):

Well you won't get a direct answer from me. That's not the policy of this website.

OpenStudy (anonymous):

then get off because you're not helpful to anyone

OpenStudy (welshfella):

If I was not helpful to anyone would I have 598 fans and over 2000 medals?

OpenStudy (anonymous):

those people are obviously nobodies who dont know what theyre doing with their life unlike me

OpenStudy (welshfella):

rubbish